# Paloma[Uncaptioned image] : A Benchmark for Evaluating Language Model Fit

### v1

@@ Abstract
Language models (LMs) commonly report perplexity on monolithic data held out from training. Implicitly or explicitly, this data is composed of domains—varying distributions of language. Rather than assuming perplexity on one distribution extrapolates to others, Perplexity Analysis for Language Model Assessment (Paloma), [^1] https://paloma.allen.ai/ measures LM fit to 585 text domains, ranging from nytimes.com to r/depression on Reddit. We invite submissions to our benchmark and organize results by comparability based on compliance with guidelines such as removal of benchmark contamination from pretraining. Submissions can also record parameter and training token count to make comparisons of Pareto efficiency for performance as a function of these measures of cost. We populate our benchmark with results from 6 baselines pretrained on popular corpora. In case studies, we demonstrate analyses that are possible with Paloma, such as finding that pretraining without data beyond Common Crawl leads to inconsistent fit to many domains.
@@ 1 Introduction
Figure 1: Perplexity macro averaged over any domains within each of the 18 top-level data sources (§2.2) in Paloma, using baselines with pretraining controls including decontamination. Evaluating on one monolithic corpus, such as C4, does not tell the complete story of model fit. Paloma lets us see when trends differ from one distribution of language to another. For instance, the 3 baselines trained on only Common Crawl data (C4, mC4-en, Falcon Refinedweb) exhibit high perplexity, sometimes with non-monotonic scaling over tokens seen, on specific evaluation sources such as The Pile, Dolma, and Dolma-100-programming-languages.
@@
Progress in AI is often catalyzed by benchmarks that define new ways of measuring progress (Deng et al., 2009, Wang et al., 2018, and Wang et al., 2019, inter alia). Language models (LMs) often report LM fit in the form of perplexity (Jelinek et al., 1977) or its logarithm, cross-entropy loss, on held out data from a model’s training distribution or a small number of traditional test sets (Chelba et al., 2013, and Merity et al., 2016, inter alia). Such measures of LM fit have been shown to improve predictably with increases in scale from more parameters, training data, and compute (Kaplan et al., 2020; Hoffmann et al., 2022). However, increasingly large training data also aggregates increasing numbers of distinct communities (Diaz & Madaio, 2023) with differing distributions of language, i.e., domains, that LMs implicitly learn to model (Aharoni & Goldberg, 2020). Does rising performance lift all data? Or do some domains capture most improvement in LM fit? How do we evaluate decisions, such as how to compose pretraining data, that determine which distributions of language are modeled? We contend that, rather than extrapolating trends from a prescriptive mix of domains, benchmarks ought to measure LM fit to many domains and inspect where fit differs.
@@
In this work we introduce Paloma, a benchmark to study LM fit on many domains. We measure perplexity on different distributions of language that we surface by sampling from 18 sources, such as C4 (Raffel et al., 2019; Dodge et al., 2021), that have metadata marking 585 textual domains, such as URL domains or academic disciplines. Beyond evaluation data, we aim to enable and enrich fair comparisons for scientific research on language modeling with the following artifacts: guidelines for experiments on LM fit, 6 baseline 1B parameter models pretrained on popular corpora, standardized inference code, and a submission process for coordinating comparable results across the research community.
@@
More than being a one-dimensional leaderboard, Paloma offers a suite of fine-grained results from submissions organized by their comparability. As reproducing pretrained models for every new project is onerous, we provide standard training controls for benchmark decontamination and training data order to orchestrate a greater density of comparisons across the research community. Submissions opt in to these, or are marked to make limitations to comparability easy to see. We also control evaluation (1) by sampling evenly from domains based on an estimate of metric variance introduced by subsampling, (2) by fixing model vocabulary where possible or otherwise using bits per byte rather than perplexity to compare, and (3) by standardizing evaluation format. Lastly, we also coordinate fair comparisons over two measures of cost, number of model parameters and training tokens, enabling assessment of hardware-agnostic Pareto efficiency and the measurement of scaling trends.
@@
In addition to curating stratified subsamples of existing datasets of fine-grained domains (Gao et al., 2020; Reid et al., 2022; Chronopoulou et al., 2022; Greenbaum & Nelson, 1996; Blodgett et al., 2016; Liang et al., 2022), we contribute new evaluation corpora constructed from held out data from the Dolma pretraining corpus (Soldaini et al., 2023) that subsample the top 100 subreddits and top 100 programming languages. Also, we repurpose corpora of fringe online communities for perplexity evaluations to measure model fit to discourse previously studied for the prevalence of toxicity and hate speech (Ribeiro et al., 2021; Zannettou et al., 2018; Papasavva et al., 2020)—an important consideration for LMs, as exposure to toxic pretraining trades off negative and positive capabilities such as toxic generation and classifying toxicity (Longpre et al., 2023). However, different lines of research will inevitably require different selections of domains beyond the scope of any one benchmark. In Paloma we focus on English and code data and aim to assemble the most fine-grained domains readily identifiable from existing metatdata, so we can begin evaluating models over stratified samples of hundreds of domains.
@@
To demonstrate possible uses of results from our benchmark, we conduct a series of case studies. We show that performance improves in almost all domains as models are scaled, but domains improve unequally. Further, across domains, perplexity is driven by strings in the vocabulary, i.e., types, that occur in most domains, but other types even get worse as models scale. Finally, our experiments isolate change in fit from which pretraining corpus is used and find that pretraining without heterogeneous data sources beyond Common Crawl leads to perplexities that do not improve consistently with number of tokens seen.
2 Paloma
Perplexity Analysis for Language Model Assessment (Paloma) is for examining LM fit to domains. We use perplexity (and related metrics; §3) to measure fit to the distributions of language represented by different domains. We take relative differences in LM fit as a proxy of model familiarity to the shared knowledge, values, and social context that position the humans producing language in a domain. While we expect contemporary LMs to have a limited fit to the most complex of these latent factors of textual domains, improving fit to all factors is important both to improve perplexity and for actual use of the LM. For example, better perplexity on a particular dialect of English suggests that model will make a better chatbot for people that speak that dialect.
Paloma comprises several types of artifacts for enabling a science of language modeling: training and evaluation guidelines for experiments on LM fit (§2.1), evaluation data for assessing fit to specific domains (§2.2), 6 pretrained baselines following training guidelines (§2.3), metrics computed by our standardized inference code conforming to our evaluation guidelines (§3), and a submission process for coordinating comparable results across the research community (§2.5).
2.1 Guidelines
We outline the principles that we adopt for assessing LM fit. To use perplexity as a meaningful measure of fit to a domain, we must account for factors in both training and evaluation that can confound results. In Table 1 we compare how previous benchmarks of language modeling have responded to these issues. We distinguish these guidelines from the controls that we use to implement these guidelines in Paloma, the technical details of which we discuss in §3.
Guideline
The Pile (Gao et al., 2020)
M2D2 (Reid et al., 2022)
C4-100-domains (Chronopoulou et al., 2022)
HELM LM Scenarios (Liang et al., 2022)
Paloma
G1 Decontamination
partial, doc-level
none
none
not required
sub-doc-level
G2 Training Order
not required
not required
not required
not required
fixed
G3 Subsampling
uniform
uniform
uniform
inherits splits
stratified
G4 Vocabulary
not required
not required
not required
not required
fixed
G5 Evaluation Format
no concat or overlap
not required
not required
API dependent
no concat or overlap
# Domains
22
216
99
14
585
Table 1: Differences between Paloma and other language modeling benchmarks with respect to guidelines (§2.1) for experiments of assessing LM fit. Paloma is the first benchmark to remove contamination across all training data, including contamination at the sub-document level. Pile only deduplicates 2 of 22 domains at document level before splitting. Paloma also fixes training data order, takes a stratified subsample of the same size from each domain based on estimated metric variance, and fixes vocabulary and evaluation format. When experiments require changes in vocabulary, bits per byte (§3) is compared instead of perplexity, following The Pile and HELM. Also following The Pile, we use an evaluation format that does not concatenate multiple documents in a single input and that uses no overlap when splitting documents longer the maximum sequence length. HELM’s inference code depends on potentially unknown inference formats used by proprietary APIs but is otherwise documented.
2.1.1 Training Guidelines
G1 Decontamination
Remove pretraining data with sub-document overlap against test data to ensure validity of perplexity evaluation.
A basic tenet of machine learning is that for test evaluation to accurately represent performance, training and test data need to be non-overlapping. However, large pretraining corpora are known to contain evaluation data and large models are known to memorize training data (Dodge et al., 2021; Elazar et al., 2023; Carlini et al., 2022). Lee et al. (2022) show in their second figure that models underestimate perplexity on evaluation documents with near duplicates in the training corpus by several points relative to models with those duplicate training documents removed. Thus benchmarks of language modeling should actively remove contaminated training data, rather than just partitioning held out splits by documents, assuming no documents overlap. The Pile applies document-level deduplication to two of their 22 domains before splitting held-out data, but its designers note that this does not prevent leakage of evaluation data more generally (Gao et al., 2020). Furthermore, spans of contaminated text within larger unrelated documents can still contribute to overestimation of performance, so decontamination should be conducted at a sub-document level. To our knowledge, Paloma is the first language modeling benchmark to require removing training data that is contaminated with respect to evaluation data.
G2 Training Order
If changes in training data order are not examined by an experiment, keep the training data order the same to control differences from recency effects.
2.1.2 Evaluation Guidelines
G3 Subsampling
Base the size of evaluation data subsamples on empirical estimates of variance over subsamples.
G4 Vocabulary
If changes in vocabulary are not examined by an experiment, keep the vocabulary the same to permit direct comparison on perplexity. If not, use bits per byte (BPB) to normalize likelihood by a segmentation intrinsic to the text.
Perplexity per token is not comparable between models with different vocabularies (Jelinek, 1998) or, by extension, different tokenizers (Mielke, 2019). Since models distribute probability over a vocabulary of tokens, models with larger vocabularies will tend to have higher perplexities than ones with smaller vocabularies. Where possible, the most rigorous solution is to impose one vocabulary on all experiments, allowing perplexity to be directly compared. Some lines of research, such as improving tokenizers, require comparisons of LM fit across vocabularies. This is possible by normalizing likelihood by a segmentation intrinsic to the text such as characters or bytes (Mielke, 2019). The Pile (Gao et al., 2020) proposes BPB (§3) as the best compromise when tokenizers are not identical, an approach we adopt as well. Paloma further establishes a standard tokenizer and vocabulary for submissions that do not need to change this experimental variable.
G5 Evaluation Format
Evaluate likelihood in a consistent format.
While perplexity is clearly defined as a function of the likelihood assigned by a model to a set of sequences, the manner in which that likelihood is computed may vary depending on how inputs are formatted for the model. The Pile (Gao et al., 2020) identify one possible variation: inferring test documents as separate inputs or concatenating them together to fill a single input. Meanwhile, Press et al. (2021) point out that documents larger than the maximum sequence length can be split either with or without overlap. We follow The Pile (Gao et al., 2020) in requiring inferences of documents in separate inputs, with documents longer than the maximum sequence length split into nonoverlapping inputs.
2.2 Evaluation data
Purpose
Source
Reference
Description
Standard language modeling benchmarks C4
Raffel et al. (2019) via Dodge et al. (2021)
Standard contemporary LM pretraining corpus automatically filtered from the April 2019 Common Crawl scrape
mC4-en
Chung et al. (2023)
The English language portion of a pretraining corpus automatically filtered from 71 Common Crawl scrapes
The Pile
Gao et al. (2020)
Standard contemporary LM benchmark from curated multi-source data including large scale non-webscraped sources
Wikitext-103
Merity et al. (2016)
A standard collection of verified “Good” and “Featured” articles on Wikipedia
Penn Treebank
Marcus et al. (1999) via Nunes (2020)
Classic Wall Street Journal benchmark with linguistic structure annotations omitted
RedPajama
Together Computer (2023)
A publicly available reproduction of the LLaMA (Touvron et al., 2023) pretraining source mixture, combining large amounts of webscraped text with smaller curated sources
Falcon Refinedweb
Penedo et al. (2023)
A corpus of English sampled from all Common Crawl scrapes until June 2023, more aggressively filtered and deduplicated than C4 and mC4-en
Dolma
Soldaini et al. (2023)
A three trillion token corpus that samples sources commonly used to train LMs in order to enable open research on pretraining data
Fine-grained domain benchmarks M2D2 S2ORC
Reid et al. (2022)
Papers from Semantic Scholar grouped by hierarchical academic field categories
M2D2 Wikipedia
Reid et al. (2022)
Wikipedia articles grouped by hierarchical categories in the Wikipedia ontology
C4-100-domains
Chronopoulou et al. (2022)
Balanced samples of the top 100 URL domains in C4 as measured by page count
Dolma-100-subreddits
Soldaini et al. (2023)
Balanced samples of the top 100 subreddits by number of posts, sourced from the Dolma Reddit subset
Dolma-100-programming-languages
Kocetkov et al. (2022) via Soldaini et al. (2023)
Balanced samples of the top 100 programming languages by number of tokens, sourced from the Dolma Stack subset
Disparities between speech communities ICE
Greenbaum & Nelson (1996) via Liang et al. (2022)
English from around the world curated by local experts, with subsets for Canada, East Africa, Hong Kong, India, Ireland, Jamaica, Philippines, Singapore, and the USA
TwitterAAE
Blodgett et al. (2016) via Liang et al. (2022)
Balanced sets of tweets classified as African American or White aligned English
Fringe sources previously studied for problematic discourse Manosphere Corpus
Ribeiro et al. (2021)
9 forums where a set of related masculinist ideologies developed over the 2000s and 2010s
Gab Corpus
Zannettou et al. (2018)
Data from 2016-2018 from an alt-right, free-speech-oriented social media platform shown to contain more hate speech than mainstream platforms
4chan Corpus
Papasavva et al. (2020)
Data from 2016-2019 from a politics subforum of an anonymity-focused forum found to contain among the highest rates of toxic content
Table 2: The 18 data sources sampled to create language modeling evaluations in Paloma. These are grouped by their purposes for inclusion (§2.2). Different lines of research will require different selections of domains; Paloma aims to enable research on differences in LM fit over the hundreds of domains that are readily available in existing metadata.
Source Validation Test Combined Domain Count Tokens per Split per Domain
C4 1,000,000 1,000,000 2,000,000 1 1,000,000
mC4-en 1,000,000 1,000,000 2,000,000 1 1,000,000
The Pile 2,199,944 2,199,333 4,399,277 22 99,984
Wikitext-103 247,969 283,134 531,103 1 265,552
Penn Treebank 89,917 101,818 191,735 1 95,868
RedPajama 699,946 700,000 1,399,946 7 99,996
Falcon Refinedweb 1,000,000 1,000,000 2,000,000 1 1,000,000
Dolma 2,999,998 2,994,903 5,994,901 6 499,575
M2D2 S2ORC 16,691,625 16,682,726 33,374,351 167 99,923
M2D2 Wikipedia 4,890,146 4,890,573 9,780,719 49 99,803
C4-100-domains 9,795,511 9,813,881 19,609,392 99 99,037
Dolma-100-subreddits 9,679,376 9,680,887 19,360,263 100 96,801
Dolma-100-programming-languages 9,999,707 9,999,906 19,999,613 100 99,998
ICE 7,290,880 7,236,065 14,526,945 17 427,263
TwitterAAE 722,905 718,358 1,441,263 2 360,316
Manosphere Corpus 1,000,000 999,915 1,999,915 9 111,106
Gab Corpus 1,000,000 1,000,000 2,000,000 1 1,000,000
4chan Corpus 1,000,000 1,000,000 2,000,000 1 1,000,000
Paloma 71,307,924 71,301,499 142,609,423 585 121,888
Table 3: Statistics of the evaluation data in Paloma. We aim for a minimum of 100 thousand tokens per domain to select a balance between inference cost and metric variance based on our empirical findings on the impact of subsampling in §3.2.1. Bold marks minimum tokens after subsampling.
In Table 2, we list the sources of evaluation data by their purposes of inclusion, and in Appendix A we detail each source individually. We show the number of tokens [^2] In this paper, token counts are always computed with the GPT-NeoX-20B tokenizer (Black et al., 2022) unless otherwise stated. and domains in each of the 18 sources in Paloma in Table 3.
In this paper, we distinguish sources from domains, although not all cases permit such easy distinction. We use source to refer to a selection of data that is characterized by the decisions of the people who curated that data, whether that curation is automatic as in scraping C4 or manual as in selecting the subcorpora of The Pile. By contrast we use domain to refer to a set of documents that belong together because they are originally produced by a group of humans that share a distinct social context. Considered as such, domains may overlap; a document’s author may belong to the set of English speakers in Jamaica and the set of AI researchers. Further note, that domains are often latent categorizations which we only approximate because complete metadata does not exist.
Also, some domains in Paloma appear in multiple sources, such as academic papers. Though The Pile and RedPajama process academic papers differently, the subcorpora on academic papers in each source represent different approximations of the same or very similar domains. However for the sake of simplicity, we make the reductive assumption of counting all 585 domains in Paloma as fully distinct.
It is beyond the scope of any one paper to prescribe an exhaustive set of domains that should be examined for a LM. Rather Paloma brings together a substantial selection of domains that are identifiable from already available metadata to demonstrate the kinds of analyses possible with hundreds of domains and rigorous experimental controls. Different research goals will motivate different definitions and selections of domains, but other researchers can apply our guidelines (§2.1) to novel fine-grained domains suitable for their research questions. One of the key advantages of evaluating a model by its fit to a collection of text representing a domain is that such domains can be identified not just by researchers who study LMs. We hope future work will identify many more domains that no one discipline would think to look at.

### v2

@@ Abstract
Evaluations of language models (LMs) commonly report perplexity on monolithic data held out from training. Implicitly or explicitly, this data is composed of domains—varying distributions of language. We introduce Perplexity Analysis for Language Model Assessment (Paloma) [^1] Dataset and links to code repository are available at https://paloma.allen.ai, a benchmark to measure LM fit to 546 English and code domains, instead of assuming perplexity on one distribution extrapolates to others. We include two new datasets of the top 100 subreddits (e.g., r/depression on Reddit) and programming languages (e.g., Java on GitHub), both sources common in contemporary LMs. With our benchmark, we release 6 baseline 1B LMs carefully controlled to provide fair comparisons about which pretraining corpus is best and code for others to apply those controls to their own experiments. Our case studies demonstrate how the fine-grained results from Paloma surface findings such as that models pretrained without data beyond Common Crawl exhibit anomalous gaps in LM fit to many domains or that loss is dominated by the most frequently occurring strings in the vocabulary.
@@ 1 Introduction
Progress in AI is catalyzed by evaluations that define new ways of measuring progress (Deng et al., 2009, Wang et al., 2018, and Wang et al., 2019, inter alia). Language models (LMs) often evaluate LM fit as loss or perplexity (Jelinek et al., 1977) on held out training data or few traditional test sets (Chelba et al., 2013; Merity et al., 2016, inter alia). These loss measures have been shown to improve predictably with increases in training compute (Kaplan et al., 2020; Hoffmann et al., 2022) and loss may predict performance on downstream tasks (Xia et al., 2022; Gadre et al., 2024; Du et al., 2024). However, scaling pretraining data aggregates more domains that LMs implicitly learn to model (Diaz and Madaio, 2023; Aharoni and Goldberg, 2020). Does rising performance lift all data? Or do some domains capture most improvement in LM fit? How do we evaluate what language distributions models learn from different pretraining data? What domains should studies evaluate loss on to measure the relationship of loss and downstream performance? To answer these questions, perplexity evaluations ought to measure LM fit to many domains, rather than extrapolating trends from a single prescriptive mix of domains.
Figure 1: Perplexity on Paloma for baselines pretrained with our experimental controls such as benchmark decontamination. We measure fit over diverse sources beyond data held-out from training. Paloma enables loss comparisons between different models, such as this figure where pretraining data is varied while all other factors are controlled. This measurement excludes documents from fringe sources and code data not supported by our decontamination approach.
@@
In this work we introduce Paloma, a benchmark to study LM fit on many domains. We measure perplexity on different distributions of language sampled from 16 sources, such as C4 (Raffel et al., 2019), that have metadata such as URLs marking 546 textual domains. Beyond evaluation data, we aim to enable and enrich fair comparisons for scientific research on language modeling with the following artifacts: guidelines for comparing LM fit, 6 baseline 1B parameter models pretrained on popular corpora, and standardized code for experiments with Paloma.
@@
As reproducing pretrained models for every new project is onerous, we provide standard training controls for benchmark decontamination and training data order to orchestrate a greater density of comparisons across the research community. We also control how Paloma is evaluated by fixing sample size per domain, model vocabulary, and inference format. Lastly, we demonstrate how to make fair comparisons over two measures of cost, number of model parameters and training tokens, enabling assessment of hardware-agnostic efficiency and the measurement of scaling trends.
@@
Among the 16 sources curated in our benchmark, we contribute two new datasets constructed from data held out of Dolma (Soldaini et al., 2024): (1) a subsample of the top 100 subreddits by number of comments, and (2) code from the top 100 programming languages by number of tokens. Also, we repurpose corpora of fringe online communities to measure LM fit to discourse previously studied for the prevalence of toxicity and hate speech (Ribeiro et al., 2021; Zannettou et al., 2018; Papasavva et al., 2020). While, capturing domains required by all possible lines of research is impossible for any one benchmark, Paloma focuses on English and code data and aims to assemble the most fine-grained domains readily identifiable from existing metadata.
@@
To demonstrate possible uses of results from our dataset, we present a series of case studies in §4. Among other findings, our experiments isolate change in fit from which pretraining corpus is used (Figure 1) and find that pretraining without heterogeneous data sources beyond Common Crawl can lead to perplexities in some domains that do not improve consistently with number of tokens seen. We also find that few vocabulary types account for most of the loss measured in perplexity.
In sum, Paloma contributes:
  1. 1.
    Curated release of the most fine-grained perplexity evaluation data in use in LM research, along with guidelines and code for standardized and rigorous perplexity evaluation.
  2. 2.
    New evaluation data for the 100 most popular subreddits and programming languages.
  3. 3.
    1B LMs pretrained on C4, mC4-en, Falcon Refinedweb, The Pile, RedPajama, and Dolma with controlled hyperparameters, token budget, benchmark decontamination, and training order for fair comparisons, along with code for others to do the same.
  4. 4.
    Case studies demonstrating analyses that are possible with Paloma, such as finding that pretraining without data beyond Common Crawl leads to inconsistent fit to many domains and that perplexity is driven by improved fit on the most common vocabulary strings.
2 Sources of evaluation data
Purpose Source Val. + Test Tokens Domains Tokens per Split per Domain
Standard language modeling benchmarks C4 (Raffel et al., 2019) 2,000,000 1 1,000,000
mC4-en (Chung et al., 2023) 2,000,000 1 1,000,000
WikiText-103 (Merity et al., 2016) 531,103 1 265,552
Penn Treebank (Marcus et al., 1999) 191,735 1 95,868
RedPajama (Together Computer, 2023) 1,399,946 7 99,996
Falcon Refinedweb (Penedo et al., 2023) 2,000,000 1 1,000,000
Dolma (Soldaini et al., 2024) 5,994,901 6 499,575
Fine-grained domain benchmarks M2D2 S2ORC (Reid et al., 2022) 33,374,351 167 99,923
M2D2 Wikipedia (Reid et al., 2022) 9,780,719 49 99,803
C4-100-domains (Chronopoulou et al., 2022) 19,609,392 99 99,037
Dolma-100-subreddits (Soldaini et al., 2024) 19,360,263 100 96,801
Dolma-100-programming-languages (Soldaini et al., 2024) 19,999,613 100 99,998
Disparities TwitterAAE (Blodgett et al., 2016) 1,441,263 2 360,316
Fringe sources Manosphere Corpus (Ribeiro et al., 2021) 1,999,915 9 111,106
Gab Corpus (Zannettou et al., 2018) 2,000,000 1 1,000,000
4chan Corpus (Papasavva et al., 2020) 2,000,000 1 1,000,000
Paloma 123,683,201 546 113,263
Table 1: The 16 data sources sampled to create language modeling evaluations in Paloma2), organized by the purpose for inclusion. These coarse-grained sources contain finer-grained domains, which use metadata to distinguish distinctive distributions of language such as a subreddit for discussing board games. Paloma aims to enable research on differences in LM fit over hundreds of domains by curating and standardizing the text datasets with the most fine-grained domains readily available from existing metadata. We target a minimum of 100 thousand tokens per domain and 1 million tokens per source to select a balance between inference cost and metric variance.
We define two terms: Sources are as existing datasets (or curated subsets there of) in use for research. Domains are fine-grained partitions of sources based on available metadata that attempt to surface a distinct and intuitive distribution of language (e.g., Wikipedia articles about visual arts or a subreddit for advice on PC builds). Paloma is derived from 16 sources further divided into 546 domains (see Table 1). [^2] Unless stated, token counts are computed with the GPT-NeoX-20B tokenizer (Black et al., 2022). Where we curate previous fine-grained corpora, we inherit their operationalization of domains, ranging from the community-driven Wikipedia ontology to expert curation and automatic classification. Where we build our own fine-grained domains from Reddit and GitHub, we make similar use of metadata about subreddits and file extensions.
Compared to monitoring monolithic validation loss during model development, interpreting LM fit to specific fine-grained domains poses unique challenges. Crucially, we must not assume better LM fit to a domain reflects improvements in the specific skills that are valued by the humans producing language in that domain (Diaz and Madaio, 2023). For instance, we might expect overlapping domains for academic papers in both Dolma and RedPajama to exhibit similar perplexities for a given model, perhaps assuming perplexity represents how much a model captures knowledge about relevant academic fields. But domains can also differ due to preprocessing when texts were collected in each source rather than from how texts were composed by their original authors. So instead of relying on LM fit to measure what we think a model should learn about a domain, we examine anomalies in domain fit to see what a model is learning. We find that the same model can have 391,171 perplexity on arXiv in RedPajama and 14 on the overlapping academic domain, peS2o, in Dolma4.1). In this approach we follow Holtzman et al. (2023) and McCoy et al. (2023) by aiming to examine model behaviors, regardless of their desirability to humans.
Also note that Paloma focuses on English and code data, as most current LMs also emphasize these types of data. However, we strongly encourage future work to explore fit to fine-grained domains in other languages.
The rest of this section addresses each source, why we include it, and how it identifies any domains it contains (all 546 domains are listed in Appendix E).